Each side of a flower garden, 30 meters long and 20 meters wide, is increased by the same amount so that the new area will be twice as big as the original area. By how many meters was each side extended?
Answer by mananth(16946)


dimensions = 30 by 20
area = 600 sq.m
let it be increased by x
increased dimensions = x+30, x+20
(x+30)(x+20)= 2*600
x^2+30x+20x+600=1200
x^2+50x+600=1200
x^2+50x-600=0
X^2+60x-10x+600=0
x(x+60)-10(x+60)=0
(x+60)(x-10)=0
x= 10 OR -60
take the positive
x= 10